18. Income statement The income statement, also known as a profit and loss (P&L) statement, provides a snapshot of a companyGreen Caterpillar Garden Supplies Inc. Income Statement for Year Ending December 31 Year 2 (Forecasted) $ Net sales Less: OpePlease give me all the answers. thanks
Answer #1
Year 1 Year 2
Net sales 150,00,000 =15000000*1.25 187,50,000
Less : Operating costs, except D&A 105,00,000 =18750000*70% 131,25,000
Less : D&A expenses 6,00,000 6,00,000
Operating income (or EBIT) 39,00,000 50,25,000
Less : Interest expenses 3,90,000 =5025000*15% 7,53,750
Pre-tax income (or EBT) 35,10,000 42,71,250
Less : Taxes (40%) 14,04,000 =4271250*40% 17,08,500
Earnings after taxes 21,06,000 25,62,750
Less : Preferred stock dividend 1,00,000 1,00,000
Earnings available to common shareholders 20,06,000 24,62,750
Less : Common stock dividend 7,37,100 8,96,963
Contribution to retained earnings 12,68,900 15,65,787
In year 2, If Green Caterpillar has 10,000 shares of preferred stock issued and outstanding, then each preferred share should except to receive annual $10 (100,000/10,000) in annual dividends.
EPS Year 1 Year 2
Shares outstanding 50,000 2006000/50000= 40.12 2462750/50000 = 49.26
EBITDA Year 1 Year 2
EBIT + D&A 3900000+600000 = 4,500,000 5025000+600000 = 5,625,000